OpenStudy (anonymous):

f(x) = 4x + 5 h(x) = √5x - 4 Find: (hf)(4) These choices 1: Square root 17 2: 5 3: 84 4:25

OpenStudy (dumbcow):

= f(4) * h(4) =(4*4+5) * sqrt(5*4-4) = 21*sqrt16 =21*4 =84